OpenStudy (anonymous):

What is the perimeter of the rectangle? Use the distributive property to rewrite the expression and solve. 2.1cm 15.3cm

OpenStudy (anonymous):

perimeter can be found using this formula: L+W+L+W

OpenStudy (anonymous):

so it would be 2.1+15.3+2.1+15.3=?

OpenStudy (anonymous):

with this in mind, can you tell me what 2.1+15.3 is?

OpenStudy (anonymous):

17.4

OpenStudy (anonymous):

correct! now what is 17.4+2.1+15.3?

OpenStudy (anonymous):

solve this and you have your answer

OpenStudy (anonymous):

p=2l+2w =2*15.3+2*2.1 =2*( +2.1) = cm @WhiteFangIII

OpenStudy (anonymous):

@wyattp17

OpenStudy (anonymous):

yes, using that, it should be easy to find out

OpenStudy (anonymous):

that is the other part i did not put it on

OpenStudy (anonymous):

omg this is history not math

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

no sweat

OpenStudy (anonymous):

its not like i will find you and burn you with fire from my maw, right? >:)